# Consumption Partial Insurance

### v3

@@ 2.1 Income process
$\sigma^{2}$ describes the dispersion of the distribution about its mean; the mean is zero by construction. $\gamma$ denotes the skewness/asymmetry of the distributions, i.e., the relative lengths of the upper and lower tails. Finally, $\kappa$ denotes the kurtosis, i.e., the tendency of the distributions to amass away from the middle $[-\sigma,\sigma]$, and thus characterizes the thickness of the tails.
We allow the central moments of shocks to depend on $t$ to reflect that different stages of the lifecycle may exhibit different amounts of dispersion, skewness, or thickness of the tails of income risk; Guvenen et al. (2021) provide some evidence for that. Otherwise, shocks across households are identically distributed within $t$. Permanent and transitory shocks are mutually independent and independent over time.
@@ 2.2 Linear consumption function
The transmission parameters vary, in principle, with $i$ (household) and $t$ (age), reflecting heterogeneity in financial wealth, among other things. The idea is that a household can use wealth to smooth consumption when income shocks hit; and the more wealth one holds the better they can self-insure. [^8] In addition, households in Blundell et al. (2016) have access to external insurance and insurance through family labor supply, while households in Theloudis (2021) differ in their preferences; both are reasons why the transmission parameters may vary with $i$ or $t$. We establish this link from wealth to $\phi_{it}^{(1)}$, $\psi_{it}^{(1)}$ in appendix A. So while the pass-through of shocks depends on wealth, it does not depend on the magnitude or sign of the shock itself. Function (4) thus implies that a household can insure big or small, good or bad shocks similarly, a point to which we return subsequently.
@@ 2.3 Quadratic consumption function
where $\Delta c_{it}$ and $\xi_{it}$ are defined as previously. The quadratic function (5) stems from a second-order Taylor approximation to the problem’s first-order conditions and the household budget constraint, as we show in appendix A. These operations and resulting expressions generalize the original approximations in BPP and in several papers thereafter.
@@
which depends on multiple parameters and on the magnitude and sign of the income shocks. As such, there is an asymmetric pass-through of good versus bad shocks into consumption. Suppose, for example, that $\phi^{(1)}>0$ and $\phi^{(2)}<0$, as our results subsequently suggest. At $v_{it}=0$ (the average transitory shock), a negative permanent shock has a higher pass-through to consumption relative to a positive shock of the same magnitude. The pass-through of negative shocks would thus be underestimated (i.e., excessive insurance concluded) if the econometrician assessed the extent of partial insurance on the basis of $\phi^{(1)}$ alone, ignoring the non-linearity in the consumption response. The size of this bias depends on the magnitude of $\phi^{(2)}$ (a more negative $\phi^{(2)}$ implies bigger bias) and on the size of the permanent shock itself (a more severe bad shock implies bigger bias). When income risk is non-Gaussian (e.g., it is left tailed), the realization of shocks can be far from their mean, thereby causing large errors in the inference of the degree of partial insurance based on $\phi^{(1)}$ alone.
A similar asymmetry exists in the transmission of good versus bad transitory income shocks. In this case, the pass-through into consumption is given by
@@ 3 Identification
Our specification of the income and consumption processes imposes restrictions on the second and higher-order moments of income and consumption that can be used to identify the various parameters. The set of parameters includes the variance ($\sigma_{\zeta_{t}}^{2}$, $\sigma_{v_{t}}^{2}$), skewness ($\gamma_{\zeta_{t}}$, $\gamma_{v_{t}}$), and kurtosis ($\kappa_{\zeta_{t}}$, $\kappa_{v_{t}}$) of the distributions of income shocks, [^11] With a slight abuse of terminology, we use the terms skewness and kurtosis to refer to the third and fourth central moments. We will be clear whenever those terms refer to standardized moments. the partial insurance parameters in the linear ($\phi^{(1)}_{t}$, $\psi^{(1)}_{t}$) and quadratic consumption functions ($\phi^{(1)}_{t}$, $\psi^{(1)}_{t}$, $\phi^{(2)}_{t}$, $\psi^{(2)}_{t}$, $\omega^{(22)}_{t}$), and the moments of the unobserved taste heterogeneity ($\sigma_{\xi_{t}}^{2}$, $\gamma_{\xi_{t}}$, $\kappa_{\xi_{t}}$). We initially restrict the partial insurance parameters to not vary across households but we subsequently explore their variation across certain demographic groups.
@@ 3.1 Income process parameters
The long sum in the first line strips income growth at $t$ of its contemporary transitory shock; its covariance with $\Delta y_{it}$ thus picks up the variance of the permanent shock. The covariance of two consecutive income growths in the second line identifies the variance of the transitory shocks due to mean reversion in the shock.
In a similar way, the skewness of shocks is identified as
@@
$\kappa_{v_{t}}$ $={\operatorname*{\mathbb{E}}}((\Delta y_{it})^{2}\times(\Delta y_{it+1})^{2})-\sigma_{\zeta_{t}}^{2}\sigma_{\zeta_{t+1}}^{2}-\sigma_{\zeta_{t}}^{2}\sum_{\tau=0}^{1}\sigma_{v_{t+\tau}}^{2}-\sigma_{\zeta_{t+1}}^{2}\sum_{\tau=-1}^{0}\sigma_{v_{t+\tau}}^{2}$
$-\sigma_{v_{t-1}}^{2}\sigma_{v_{t}}^{2}-\sigma_{v_{t-1}}^{2}\sigma_{v_{t+1}}^{2}-\sigma_{v_{t}}^{2}\sigma_{v_{t+1}}^{2}.$
@@ 3.3 Quadratic consumption function parameters
This depends not only on the variance of permanent and transitory income shocks, as in the linear case, but also on their skewness and kurtosis. Contrary to the linear case, however, the higher-order moments of consumption growth depend on higher than fourth-order moments of the shocks, which we do not model.
@@
where square matrix $\mathbf{A}$ depends exclusively on the second, third, and fourth moments of income shocks. [^14] The matrix of coefficients is given by $\mathbf{A}=\left(\begin{array}[]{lllll}\sigma_{\zeta_{t}}^{2}&\sigma_{v_{t}}^{2}&\gamma_{\zeta_{t}}&\gamma_{v_{t}}&0\\ \gamma_{\zeta_{t}}&\gamma_{v_{t}}&\kappa_{\zeta_{t}}+\sigma_{\zeta_{t}}^{2}(\sigma_{v_{t}}^{2}+\sigma_{v_{t-1}}^{2})&\kappa_{v_{t}}+\sigma_{v_{t}}^{2}(\sigma_{\zeta_{t}}^{2}+\sigma_{v_{t-1}}^{2})&2\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}\\ 0&-\sigma_{v_{t}}^{2}&0&-\gamma_{v_{t}}&0\\ 0&\gamma_{v_{t}}&\sigma_{\zeta_{t}}^{2}(\sigma_{\zeta_{t+1}}^{2}+\sigma_{v_{t}}^{2}+\sigma_{v_{t+1}}^{2})&\kappa_{v_{t}}+\sigma_{v_{t}}^{2}(\sigma_{\zeta_{t+1}}^{2}+\sigma_{v_{t+1}}^{2})&0\\ 0&-\gamma_{v_{t}}&-\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}&-\kappa_{v_{t}}&-\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}\\ \end{array}\right).$ , [^15] The variance of consumption taste heterogeneity is identified from ${\operatorname*{\mathbb{E}}}((\Delta c_{it})^{2})$. Its skewness and kurtosis enter the third and fourth moments of consumption growth. Those depend on higher than fourth-order moments of income, which we do not model, so they are generally not identified in the current setting. We establish in appendix B that matrix $\mathbf{A}$ is full rank and invertible for any distribution of income shocks except Bernoulli (for which there is no empirical evidence anyway). [^16] If the distributions of both shocks are Bernoulli, then there exist some parameter configurations for which matrix $\mathbf{A}$ is not full rank. Assuming away Bernoulli distributions is thus an identifying assumption. The matrix is invertible even if the distributions of shocks are normal and have the same variance between them. In the nonlinear specification, tail shocks are associated with a potentially different pass-through from that of average shocks; but identification of such pass-through does not require those shocks to be non-Gaussian.
@@
As seen in (8) and matrix $\mathbf{A}$, the transmission parameters can no longer be identified with second-order income moments alone. Higher-order moments of income play an essential role in identification, which is intuitive: we want to identify the impact of tail shocks, so we must use information on said tails. By contrast, no higher-order moments of consumption are needed; the identifying moments in (8) are covariances between consumption growth and various powers of income growth. This is an advantage over methods that require observing the full consumption distribution (e.g., Arellano et al., 2017, need the entire consumption-income joint distribution) because, while in principle we may estimate higher moments of consumption growth in the PSID, we cannot be certain how reliable these estimates are. This is in contrast to higher income moments, for which Guvenen et al. (2021) is a natural benchmark given their use of administrative data. We return to this point in section 4.
@@
also failing to identify the true pass-through of transitory shocks in the presence of nonlinear transmission, i.e., expression (7). If $\gamma_{v_{t}}<0$ and $\psi_{t}^{(2)}>0$, as our baseline estimates suggest, linearity biases $\psi_{t}^{(1)}$ downwards. This helps explain the discrepancy between BPP (who find little pass-through of transitory shocks) and studies based on natural experiments (who find large consumption responses to said shocks). [^17] The discrepancy is between studies that use covariance restrictions in the joint distribution of income and consumption (they find small pass-through of transitory shocks) and studies that employ natural experiments (e.g., tax refunds or lotteries; they find large pass-through). Commault (2022) explains the discrepancy on the basis of neglected effects from past transitory shocks on consumption, while Crawley and Kuchler (2023) explain it on the basis of neglected time aggregation in the data. Both are alternative types of misspecification in the linear consumption function/empirical implementation vis-à-vis our focus on nonlinearities here.
@@ 4.1 Data
As is evident from the previous section, panel data on income and consumption are needed to identify the moments of the permanent and transitory shocks and the degrees of insurance to them. The Panel Study of Income Dynamics (PSID) provides such data for a representative and continuously evolving sample of U.S. households since $1968$. For most of this period, the PSID collected consumption expenditures only on food items (food representing a declining fraction of the consumer’s basket over time). Since $1999$, however, the PSID was drastically redesigned in that it started collecting information over a broader set of consumption categories, currently comprising over 70% of consumption in the National Income and Product Accounts (Blundell et al., 2016). We, therefore, use the PSID over the period between $1999$ and $2019$, the last available wave when writing this paper; the data are available only biennially. This is in contrast to BPP, who use PSID data over 19801980198019801992199219921992 and rely on consumption imputations from the Consumer Expenditure Survey (CEX).
@@
Our baseline sample consists of married couples in the PSID with a male spouse aged $30$ to $65$. We focus on a sample that is representative of the U.S. population, eliminating the supplementary low-income subsample (SEO). Households are also removed in the event of a break-up or remarriage; our analysis thus focuses on income risk rather than on divorce or other household dissolution factors. Naturally, we require non-missing data on income, expenditure, and basic demographics. [^18] We also remove observations with extreme jumps in income or consumption in consecutive years (jump: $\geq$10-fold increase, or $\geq$90reduction, or the bottom 0.25% of the distribution of products of growth rates between $t$ and $t-1$), i.e., large increases followed by large drops, likely reflecting measurement error. Our final sample consists of 20,866 household-year observations; on average, a household is observed for $5.6$ periods, which corresponds to about $10$ calendar years given the biennial nature of the data.
Table 1: Descriptive statistics
Baseline sample Low wealth High wealth No college Some college
Mean Median Mean Mean Mean Mean
(1) (2) (3) (4) (5) (6)
Consumption
Nondurable+services 26,766 22,895 23,106 29,852 22,137 28,908
Food at home 8,256 7,609 7,861 8,580 7,462 8,630
Utilities 3,515 3,238 3,270 3,721 3,394 3,571
Food out 3,318 2,511 2,683 3,840 2,649 3,634
Health 1,614 863 1,390 1,803 1,429 1,699
Transportation 5,866 4,900 5,398 6,261 5,436 6,065
Education 3,202 0 1,529 4,614 1,262 4,100
Childcare 977 0 937 1,009 543 1,181
Rent (or rent eq.) 18,105 13,821 10,949 23,991 11,647 21,149
Home insurance 894 747 654 1,092 702 985
Income
Disposable income 142,910 113,581 98,527 179,416 97,675 164,239
Taxable+transfer inc. 137,277 109,003 93,358 173,401 90,810 159,187
Earnings 118,423 97,154 87,005 144,266 78,565 137,217
Male 81,857 62,185 56,817 102,452 50,199 96,783
Female 36,567 29,183 30,187 41,814 28,366 40,433
Wealth 547,761 191,193 69,731 950,823 284,036 669,788
Demographics
% working (male) 0.89 1.00 0.92 0.88 0.86 0.91
% working (female) 0.79 1.00 0.81 0.78 0.79 0.79
Age (male) 46 45 43 49 46 46
% some college (male) 0.68 1.00 0.55 0.79 0.00 1.00
% some college (female) 0.71 1.00 0.62 0.78 0.44 0.84
Periods per household 5.63 5.00 4.89 6.45 5.21 5.86
Observations 20,866 9,417 11,449 6,686 14,180
@@
Table 1 presents summary statistics over the sample period 19991999199919992019201920192019. About $89\%$ of men and $79\%$ of women work for pay (but some households have intermittent employment); $68\%$ of men and $71\%$ of women have had some college education. Our measure of income is household disposable income, namely the sum of taxable income (e.g., earnings and asset income), transfers, and social security income of the couple and of other family members, minus taxes paid. Average disposable income is \$142,910 in 2018 prices. [^19] We express all monetary figures in 2018 prices because earnings are 1-year retrospective and our latest data are from 2019. We deflate using the Consumer Price Index from the Bureau of Labor Statistics. Our measure of consumption is real expenditure on nondurable goods and services, comprising food (at home and outside), utilities, out-of-pocket health expenses, public transport, vehicle expenses (gasoline, parking, insurance), education, and child care. Average real consumption is \$26,766, with food at home accounting for about $30\%$ of this. Housing, including rent and home insurance, is the main expenditure we exclude; we do so for comparability with BPP who also exclude it. Including housing does not change any of the subsequent results meaningfully. Finally, table 1 reports summary statistics also for samples sorted by wealth and education; we postpone the discussion of those samples to section 5.4.

### v4

@@ 2.1 Income process
$\sigma^{2}$ describes the dispersion of the distribution about its mean; the mean is zero by construction. $\gamma$ denotes the skewness/asymmetry of the distribution, i.e., the relative length of the upper and lower tails. Finally, $\kappa$ denotes the kurtosis, i.e., the tendency of the distribution to amass away from the middle $[-\sigma,\sigma]$, and thus characterizes the thickness of the tails.
We allow the central moments of shocks to depend on $t$ to reflect that different stages of the lifecycle may exhibit different amounts of dispersion, skewness, or thickness of the tails of income risk; Guvenen et al. (2021) provide some evidence for that. Otherwise, shocks across households are identically distributed given $t$. Permanent and transitory shocks are mutually independent and independent over time.
@@ 2.2 Linear consumption function
The transmission parameters vary, in principle, with $i$ (household) and $t$ (age), reflecting heterogeneity in financial wealth, among other things. [^8] In addition, households in Blundell et al. (2016) differ in their access to external insurance and insurance through family labor supply, while households in Theloudis (2021) differ in their preferences. The idea is that a household can use wealth to smooth consumption when income shocks hit; and the more wealth one holds the better they can self-insure. We establish this link from wealth to $\phi_{it}^{(1)}$, $\psi_{it}^{(1)}$ in appendix A. So while the pass-through of shocks depends on wealth, it does not depend on the magnitude or sign of the shock itself. Function (4) thus implies that a household can insure big or small, good or bad shocks similarly, a point to which we return subsequently.
@@ 2.3 Quadratic consumption function
where $\Delta c_{it}$ and $\xi_{it}$ are defined as previously. The quadratic function (5) stems from a second-order Taylor approximation to the problem’s first-order conditions and the household budget constraint, as we show in appendix A. These operations and resulting expressions generalize the original approximations in BPP and in several papers thereafter. [^10] We do not go beyond a quadratic approximation for reasons of tractability and because higher-order terms in a regression of $\Delta c_{it}$ on a polynomial in $\Delta y_{it}$ are imprecisely estimated.
@@
which depends on multiple parameters and on the magnitude and sign of the income shocks. As such, there is an asymmetric pass-through of good versus bad shocks into consumption. Suppose, for example, that $\phi^{(1)}>0$ and $\phi^{(2)}<0$, as our results subsequently suggest. At $v_{it}=0$ (the average transitory shock), a negative permanent shock has a higher pass-through to consumption relative to a positive shock of the same magnitude. The pass-through of negative shocks would thus be underestimated (i.e., excessive insurance concluded) if the econometrician assessed the extent of partial insurance on the basis of $\phi^{(1)}$ alone, ignoring the non-linearity in the consumption response. The size of this bias increases with the magnitude of $\phi^{(2)}$ and the size of the permanent shock itself. When income risk is non-Gaussian (e.g., it is left tailed), the realization of shocks can be far from their mean, thereby causing large errors in the inference of the degree of partial insurance based on $\phi^{(1)}$ alone.
A similar asymmetry exists in the transmission of good versus bad transitory shocks. In this case, the pass-through into consumption is given by
@@ 3 Identification
Our specification of the income and consumption processes imposes restrictions on the second and higher-order moments of income and consumption that can be used to identify the various parameters. The set of parameters includes the variance ($\sigma_{\zeta_{t}}^{2}$, $\sigma_{v_{t}}^{2}$), skewness ($\gamma_{\zeta_{t}}$, $\gamma_{v_{t}}$), and kurtosis ($\kappa_{\zeta_{t}}$, $\kappa_{v_{t}}$) of the distributions of income shocks, [^12] With a slight abuse of terminology, we use the terms skewness and kurtosis to refer to the third and fourth central moments. We will be clear whenever those terms refer to standardized moments. the partial insurance parameters in the linear ($\phi^{(1)}_{t}$, $\psi^{(1)}_{t}$) and quadratic consumption functions ($\phi^{(1)}_{t}$, $\psi^{(1)}_{t}$, $\phi^{(2)}_{t}$, $\psi^{(2)}_{t}$, $\omega^{(22)}_{t}$), and the moments of the unobserved taste heterogeneity ($\sigma_{\xi_{t}}^{2}$, $\gamma_{\xi_{t}}$, $\kappa_{\xi_{t}}$). We initially restrict the partial insurance parameters to not vary across households but we subsequently explore their variation across certain demographic groups and by the household’s position in the distribution of income.
@@ 3.1 Income process parameters
The long sum in the first line strips income growth at $t$ of its contemporary transitory shock; its covariance with $\Delta y_{it}$ thus picks up the variance of the permanent shock. The covariance of two consecutive income growths in the second line identifies the variance of the transitory shocks due to mean reversion in the shock. In a similar way, skewness is identified from
@@
$\kappa_{v_{t}}$ $={\operatorname*{\mathbb{E}}}((\Delta y_{it})^{2}\times(\Delta y_{it+1})^{2})-\sigma_{\zeta_{t}}^{2}\sum_{\tau=0}^{1}\sigma_{v_{t+\tau}}^{2}-\sigma_{\zeta_{t+1}}^{2}\sum_{\tau=-1}^{0}\sigma_{v_{t+\tau}}^{2}-\sigma_{v_{t-1}}^{2}\sum_{\tau=0}^{1}\sigma_{v_{t+\tau}}^{2}$
$-\sigma_{\zeta_{t}}^{2}\sigma_{\zeta_{t+1}}^{2}-\sigma_{v_{t}}^{2}\sigma_{v_{t+1}}^{2}.$
@@ 3.3 Quadratic consumption function parameters
This depends not only on the variance of permanent and transitory income shocks, as in the linear case, but also on their skewness and kurtosis. Contrary to the linear case, however, the higher-order moments ${\operatorname*{\mathbb{E}}}((\Delta c_{it})^{3})$ and ${\operatorname*{\mathbb{E}}}((\Delta c_{it})^{4})$ depend on higher than fourth-order moments of the shocks, which we do not model.
@@
where square matrix $\mathbf{A}$ depends exclusively on the second, third, and fourth moments of income shocks. [^15] The matrix of coefficients is given by $\mathbf{A}=\left(\begin{array}[]{lllll}\sigma_{\zeta_{t}}^{2}&\sigma_{v_{t}}^{2}&\gamma_{\zeta_{t}}&\gamma_{v_{t}}&0\\ \gamma_{\zeta_{t}}&\gamma_{v_{t}}&\kappa_{\zeta_{t}}+\sigma_{\zeta_{t}}^{2}(\sigma_{v_{t}}^{2}+\sigma_{v_{t-1}}^{2})&\kappa_{v_{t}}+\sigma_{v_{t}}^{2}(\sigma_{\zeta_{t}}^{2}+\sigma_{v_{t-1}}^{2})&2\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}\\ 0&-\sigma_{v_{t}}^{2}&0&-\gamma_{v_{t}}&0\\ 0&\gamma_{v_{t}}&\sigma_{\zeta_{t}}^{2}(\sigma_{\zeta_{t+1}}^{2}+\sigma_{v_{t}}^{2}+\sigma_{v_{t+1}}^{2})&\kappa_{v_{t}}+\sigma_{v_{t}}^{2}(\sigma_{\zeta_{t+1}}^{2}+\sigma_{v_{t+1}}^{2})&0\\ 0&-\gamma_{v_{t}}&-\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}&-\kappa_{v_{t}}&-\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}\\ \end{array}\right).$ , [^16] The variance of consumption taste heterogeneity is identified from ${\operatorname*{\mathbb{E}}}((\Delta c_{it})^{2})$. Its skewness and kurtosis enter the third and fourth moments of consumption growth. Those depend on higher than fourth-order moments of income, which we do not model, so they are generally not identified in the current setting. We establish in appendix B that matrix $\mathbf{A}$ is full rank and invertible for any distribution of income shocks except Bernoulli (for which there is no empirical evidence anyway). [^17] If the distributions of both shocks are Bernoulli, then there exist some parameter configurations for which matrix $\mathbf{A}$ is not full rank. Assuming away Bernoulli distributions is thus an identifying assumption. Identification of the parameters does not require those shocks to be non-Gaussian; the matrix is invertible even if the distributions of shocks are normal and have the same variance between them (i.e., $\sigma_{\zeta_{t}}^{2}=\sigma_{v_{t}}^{2}$).
In the nonlinear specification, tail shocks are associated with a different pass-through from that of average shocks, regardless of the precise nature of the underlying distribution of shocks. The transmission parameters, however, can no longer be identified with second-order income moments alone – inspect matrix $\mathbf{A}$ in (8). Skewness and kurtosis, both defined also in the Gaussian case, are crucial for identifying the pass-through of such shocks, which is intuitive: we want to identify the impact of tail shocks, so we must use information on said tails. In appendix B, we investigate what different levels of income skewness and kurtosis imply for the transmission parameters.
@@
Contrary to income, no higher-order moments of consumption are needed; the identifying moments in (8) are covariances between consumption growth and various powers of income growth. This is an advantage over quantile methods that require observing the entire consumption distribution, e.g., ABB, because, while we may estimate higher moments of consumption growth in the PSID, there is no established population benchmark to assess them against. This is in contrast to higher income moments, for which Guvenen et al. (2021) is a natural benchmark given their use of administrative population data.
@@
also failing to identify the true pass-through of transitory shocks in the presence of nonlinear transmission, i.e., expression (7). If $\gamma_{v_{t}}<0$ and $\psi_{t}^{(2)}>0$, as our baseline estimates suggest, linearity biases $\psi_{t}^{(1)}$ downwards. This may help bridge the discrepancy between covariance based methods (who find little pass-through of transitory shocks) and studies based on natural experiments (who usually find large consumption responses to said shocks).
Our estimation procedure, which we describe in the next section, is linked transparently to our identification argument. Effectively, we bring (8) to the data, along with some over-identifying restrictions, using moments estimation.
@@ 4.1 Data
As is evident from the previous section, panel data on income and consumption are needed to identify the moments of the permanent and transitory shocks and the degrees of insurance to them. The Panel Study of Income Dynamics (PSID) provides such data for a representative and continuously evolving sample of U.S. households since $1968$. For most of this period, the PSID collected expenditures only on food items. Since $1999$, however, the PSID was drastically redesigned in that it started collecting information over a broader set of consumption categories, currently comprising over 70% of consumption in the National Income and Product Accounts (Blundell et al., 2016). We, therefore, use the PSID over the period between $1999$ and $2019$, the last available wave when we started working on this paper; the data are available only biennially. This is in contrast to BPP, who use PSID data over 19801980198019801992199219921992 and rely on consumption imputations from the Consumer Expenditure Survey (CEX).
@@
Our baseline sample consists of married couples in the PSID with a male spouse aged $30$ to $65$. We focus on a sample that is representative of the U.S. population, eliminating the supplementary low-income subsample. Households are also removed in the event of a break-up or remarriage; our analysis thus focuses on income risk rather than on divorce or other household dissolution factors. Naturally, we require non-missing data on income, expenditure, and basic demographics. [^18] We also remove observations in the bottom 0.25% of the distributions of $\Delta y_{it}\Delta y_{it+1}$ and $\Delta c_{it}\Delta c_{it+1}$, i.e., large increases followed by large drops, likely reflecting measurement error, and observations with extreme jumps in income or consumption (jump: $\geq$10-fold increase or $\geq$90reduction), which likely also reflect measurement error. Our results are not sensitive to this sample selection choice. Our final sample consists of 20,866 household-year observations; on average, a household is observed for $5.6$ periods, which corresponds to about $10$ calendar years given the biennial nature of the data.
@@
Appendix table C.1 presents summary statistics over the sample period 19991999199919992019201920192019. About $89\%$ of men and $79\%$ of women work for pay (but some households have intermittent employment); $68\%$ of men and $71\%$ of women have had some college education. Our measure of income is household disposable income, namely the sum of taxable income (e.g., earnings and asset income), transfers, and social security income of the couple and of other family members, minus taxes paid. Average disposable income is \$142,910 in 2018 prices. [^19] We express all monetary figures in 2018 prices because earnings are 1-year retrospective and our latest data are from 2019. We deflate using the Consumer Price Index from the Bureau of Labor Statistics. Our measure of consumption is real expenditure on nondurable goods and services, comprising food (at home and outside), utilities, out-of-pocket health expenses, public transport, vehicle expenses (gasoline, parking, insurance), education, and child care. Average real consumption is \$26,766, with food at home accounting for about $30\%$ of this. Housing, including rent and home insurance, is the main expenditure we exclude; we do so for comparability with BPP who also exclude it. Including housing does not meaningfully change the subsequent results.